OpenStudy (anonymous):

dannae surveys her class to find out thier favrite food 3/6 of the students chose pizza and 2/8 of the students chose hamburagas what fraction student chose pizza either hamburargas dannae surveys her class to find out thier favrite food 3/6 of the students chose pizza and 2/8 of the students chose hamburagas what fraction student chose pizza either hamburargas @Mathematics

OpenStudy (anonymous):

3/6+2/8= 12/24+6/24=18/24=3/4

OpenStudy (anonymous):

3/6 + 2/8 equals 3/6 x 4 =12/24 plus 2/8 x 3 =6/24 add both 18/24 equals 3/4